1/4

Christian Morgenstern, Christian Morgenstern, Christian Morgen- stern, Christian Morgenstern, Christian Morgenstern,

The Bans- hee  (An Approach) Der Wer- wolf El Hechi- cero Varulven The Werewolf

translated by Max Knight translated by Hans Eiseneck translated by Sven Collberg translated by Alexander Gross

One night, a bans-

hee slunk away from mate and

child, and in the gloom

went to a village teacher’s tomb,

requesting him: “Inflect me, pray.”

Ein Werwolf eines Nachts entwich

von Weib und Kind und sich begab

an eines Dorfschullehrers Grab

und bat ihn: �Bitte, beuge mich!“ Un hechicero, con esposa vino un día a la fosa de su maestro, y al profesor pidió: “¿Me cuentas? ¡Por favor!”

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

A Werewolf, troubled by his name, Left wife and brood one night and came

To a hidden graveyard to enlist The aid of a longdead philologist.

The village teacher climbed up straight

upon his grave stone with its plate

and to the appari-

tion said who meekly knelt

before the dead:

Der Dorfschulmeister stieg hinauf

auf seines Blechschilds Messingknauf

und sprach zum Wolf, der seine Pfoten

geduldig kreuzte vor dem Toten:

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero: mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

“Oh sage, wake up, please don’t berate me,”

He howled sadly, “Just conjugate me.”

The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

…  …  …  …  …

2/4

“The banSHEE, in the subject’s place;

the banHERS, the possessive case.

The banHER, next, is what they call

objective case– and that is all.”

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Genitiv sodann,

dem Werwolf, Dativ, wie man’s nennt,

den Werwolf, - damit hat’s ein End!“ “El hechiuno” le explica, “el hechidos” después indica, “el hechitres” dice al fin, “con eso basta ¡chiquitín!”

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det varulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense,

And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

The banshee marveled at the cases

and writhed with pleasure, making faces,

but said: “You did

not add, so far, the plural to the

singular!”

Dem Werwolf schmeichelten die Fälle

er rollte seine

Augenbälle. �Indessen“, bat

er, �füge doch zur Einzahl auch

die Mehrzahl noch!“ El hechicero, conmovido por lo que capta su oído, al maestro ruega, encantado, que le adjunte un quebrado.

Varulven myste gladåt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused– again he tried:

The teacher, though, admitted then

that this was not

within his ken. “While bans are

frequent”, he advised,

“A she cannot be plurized.”

Der Dorfschullehrer aber musste

gestehn, dass er von ihr nichts wusste.

Zwar Wölfe gab’s in großer Schar,

doch �Wer“ gab’s nur im Sinular.

Pero el profesor confiesa que nada sabe con certeza: “Hay muchos hechos. Hechiceros

únicamente hay enteros.”

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

…  …  …  …  …

3/4

The banshee, ri-

sing clammily, wailed: “What

about my family?”

Then, being not a learned creature,

said humbly “Thanks” and left the teacher.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

El brujo se volvió muy triste: “Mi cara mitad, ¿no la viste?” Ya que no era sabio con gratitud se despidió.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste existera.

The Werewolf knew better–his sons still slept At home, and homewards now he crept,

Happy, humble, without apology For such folly of phi-

lology.

…  …  …  …  …

4/4

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_ poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into hispana by Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Swedish by Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu  ĉi  angli-
go  troviĝas  en

http: // bertilow. com/ literaturo/ lupfantomo. html .